How do I solve for substitution?
x=3y-14
x=4y-9

Answer by nyc_function(2741)

x=3y-14
x=4y-9
Look carefully.
As you can see, x is given in terms of y for both equations.
Simply equate both equations and solve for y.
3y - 14 = 4y - 9
3y - 4y = 14 - 9
-y = 5
y = 5/-1
y = -5
Now, plug this value of y just found into either original equations given to solve for x.
I will use x = 3y - 14 but you can use the other equation if you like.
x = 3(-5) - 14
x = -15 - 14
x = -29
The solution is the point (-29, -5).
In other words, x = -29 and y = -5.
Both equations meet at the point (-29, -5) on the coordinate plane.